# ARR: Question Answering with Large Language Models via Analyzing, Retrieving, and Reasoning

### v1

@@ 4.1 Datasets
Table 2: Main experiments. The zero-shot performance (Accuracy %) of the LLaMA3-8B-Chat model on various multiple-choice QA datasets using different answer trigger sentences $\phi$. (1) Baseline: $\phi$ is “Answer:”; (2) CoT Kojima et al. (2022): $\phi$ is “Answer: Let’s think step by step.”; (3) ARR: our method that elicits intent analyzing, information retrieving, and step-by-step reasoning.
@@ 5 Main Experiments
In this section, we present the main experimental results, which verify our hypothesis by demonstrating the effectiveness of the proposed ARR method.
@@ 5.1 QA Performance
The main experiments test the zero-shot QA performance of LLaMA3-8B-Chat Dubey et al. (2024) on various multiple-choice QA datasets. We compare three methods (i.e., Baseline, zero-shot CoT Kojima et al. (2022), and ARR), and the only difference between them is the answer trigger sentence $\phi$ shown in Figure 2. The results in Table 2 demonstrate that our ARR method boosts the Baseline method by a large margin, with an improvement of $+4.1\%$ on average. Moreover, ARR consistently outperforms zero-shot CoT prompting across all QA datasets, highlighting its universal superiority in various task types including reading comprehension, commonsense reasoning, world knowledge, and multitask understanding.
@@ 5.2 Ablation Study
We observe that all the single-component ARR settings (➁, ➂, and ➃) outperform the Baseline method (➄) by a large margin, which verifies that each ARR component contributes positively. In addition, ➃ outperforms the CoT method in Table 2, which means the “Reasoning” prompting of our ARR method (i.e., “answer the question with step-by-step reasoning”) is better than zero-shot CoT prompting (i.e., “think step by step”) for QA tasks. Furthermore, the complete ARR method (➀) has a higher accuracy score than the Retrieving-only (➂) and Reasoning-only (➃) methods, meaning the intent Analyzing benefits the other two “R” parts.

### v2

@@ 4.1 Datasets
Table 2: Main experiments. The zero-shot performance (Accuracy %) of the LLaMA3-8B-Chat model on various multiple-choice QA datasets using different answer trigger sentences $\phi$. (1) w/o Reason: directly selecting an option without rationales; (2) Baseline: $\phi$ is “Answer:”; (3) CoT Kojima et al. (2022): $\phi$ is “Answer: Let’s think step by step.”; (4) ARR: our method that elicits intent analysis, information retrieval, and step-by-step reasoning.
@@ 5.1 QA Performance
The main experiments test the zero-shot QA performance of LLaMA3-8B-Chat Dubey et al. (2024) on various multiple-choice QA datasets. The only difference between Baseline, zero-shot CoT Kojima et al. (2022), and ARR is the answer trigger sentence $\phi$ shown in Figure 2. The results in Table 2 demonstrate that our ARR method boosts the Baseline method by a large margin, with an improvement of $+4.1\%$ on average. In addition, ARR consistently outperforms zero-shot CoT prompting across all QA datasets, highlighting its universal superiority in various task types including reading comprehension, commonsense reasoning, world knowledge, and multitask understanding. Moreover, the “w/o Reason” method, which directly selects options without relying on rationales ($r_{i}$ in Eq. 2), performs significantly worse, emphasizing the benefits of our two-stage QA approach.
@@ 5.2 Ablation Study
We observe that all the single-component ARR settings (➁, ➂, and ➃) outperform the Baseline method (➄) by a large margin, which verifies that each ARR component contributes positively. In addition, ➃ outperforms the CoT method in Table 2, which means the “Reasoning” prompting of our ARR method (i.e., “answer the question with step-by-step reasoning”) is better than zero-shot CoT prompting (i.e., “think step by step”) for QA tasks. Furthermore, the complete ARR method (➀) has a higher accuracy score than the Retrieving-only (➂) and Reasoning-only (➃) methods, meaning the intent analysis benefits the other two “R” parts.